OpenStudy (anonymous):

I need help finding the derivative of x(x+4)

OpenStudy (anonymous):

Well lets use distributive property \[x(x+4)=x^2+4x\] \[ \frac{d}{dx}{x^2}+\frac{d}{dx}4x\]

OpenStudy (anonymous):

Can you solve from here?

OpenStudy (anonymous):

All you gotta use is the power rule which states \[ \frac{d}{dx}x^b = b*x^{b-1} \]

OpenStudy (anonymous):

No. But I do see what you did. Do you mind helping me with the rest? Please and thank you.

OpenStudy (anonymous):

For example \[ \frac{d}{dx}x^4=4x^{4-1}=4x^3\] Another Example which I will show in many steps so you can follow. \[ \frac{d}{dx} 5x^2=5(x^2)=5(2x^{2-1})=5(2x)=10x\]

OpenStudy (anonymous):

@Tyese would you like to give it a try?

OpenStudy (anonymous):

I'm trying now. It's just confusing with the + sign

OpenStudy (anonymous):

Well Just do each one separately and then add them together at the end

OpenStudy (anonymous):

Ok. Give me one sec

OpenStudy (anonymous):

Umm you can show me what u have and I can guide you

OpenStudy (anonymous):

Ok I got 2x

OpenStudy (anonymous):

Yes :D

OpenStudy (anonymous):

That's right????

OpenStudy (anonymous):

YUP :) Now you gotta derive the second term

OpenStudy (anonymous):

Ok. Trying now

OpenStudy (anonymous):

Ok I'm kind of stuck on this one. Is it 1.?

OpenStudy (anonymous):

Well now its \[ \frac{d}{dx}4x=4\frac{d}{dx}(x^{1})=4(1*x^{1-1})=4(1*1)=4\] Umm 2 things \[x=x^1\] and \[x^0=1\]

OpenStudy (anonymous):

Do you follow?

OpenStudy (anonymous):

Ok. So the 4 is what threw me off.

OpenStudy (anonymous):

So whats ur answer now. Lets pull it alltogether

OpenStudy (anonymous):

8x

OpenStudy (anonymous):

hmmm how did you get that? Can you show me?

OpenStudy (anonymous):

Yes. I did 2x * 4

OpenStudy (anonymous):

close. Remember there is a plus sign not a multiplication sign between the 2 derivatives

OpenStudy (anonymous):

Oh. So the answer is 2x+4

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

Do you have any other questions? Would you like to solve anther one together?

OpenStudy (anonymous):

Yay!!! Yes

OpenStudy (anonymous):

I'm grabbing my book

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

That was my last one!!!! I can't thank you enough. Are you always available?

OpenStudy (anonymous):

I come on randomly. Whenever I am in the mood of helping

OpenStudy (anonymous):

But you can always tag me and if I am around I will help

OpenStudy (anonymous):

Thank you so much!!! I may pass my quiz now

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Would the answer be different if I had a square root?

OpenStudy (anonymous):

Definitely

OpenStudy (anonymous):

Can we try it witha square root?

OpenStudy (anonymous):

\[\sqrt{3x} =\large x^{\frac{1}{3}}\]

OpenStudy (anonymous):

Basically in general terms \[\sqrt{x}= x^{\frac{1}{2}}\]

OpenStudy (anonymous):

Can you write out the complete equation so that I dont mistaken the equation

OpenStudy (anonymous):

Can we do the same problem with a square root? Likethe derivative of the square root of x(x+4)

OpenStudy (anonymous):

\[\frac {d}{dx}\sqrt{x(x+4)}\]

OpenStudy (anonymous):

hmmm do you know the chain rule?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Umm are you supposed to know it?

OpenStudy (anonymous):

Yes:-(.

OpenStudy (anonymous):

Okkk let me try to explain it. Let me think abt it one sec

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So I will try to simply explain it but its gonna be hard.

OpenStudy (anonymous):

Ok, I'm ready

OpenStudy (anonymous):

First we will solve the outer most bracket and then we will go inwards. I will explain using the example it may be easiest

OpenStudy (anonymous):

So lets take our \[\frac{d}{dx}\sqrt{x(x+4)}=\frac{d}{dx}\sqrt{x^2+4x}\] Ok so far so good? I am just simplifying it

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\frac{d}{dx} \sqrt{x^2+4x}=\frac{d}{dx}(x^2+4)^{\large\frac{1}{2}}\]

OpenStudy (anonymous):

Do you follow that?

OpenStudy (anonymous):

I'm still here. Just so you know...

OpenStudy (anonymous):

Ohhhh I can access this

OpenStudy (anonymous):

OMG Open Study is totally misbehaving.

OpenStudy (anonymous):

So I found some chain rule nots!!! I'm trying this too

OpenStudy (anonymous):

Ok but did you follow what I did so far? I have not used the chain rule yet. But I will if you follow my simplification

OpenStudy (anonymous):

I Are you there??? I think I have something.

OpenStudy (anonymous):

I am here :) Show me what you have lol

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Are you gone?

OpenStudy (anonymous):

\[ \frac{d}{dx} (x^2+4x)^{\large \frac{1}{2}}\] So first we treat whatever is in the bracket like a variable and differentiate So we would be to use the power rule \[ \frac{1}{2}(x^2+4x)^{\large \frac{1}{2}-1}=\frac{1}{2}(x^2+4x)^{\large -\frac{1}{2}}\] Then the next step would be to differentiate what is inside the bracket. \[\frac{1}{2}(x^2+4x)^{\large -\frac{1}{2}}* \frac{d}{dx}(x^2+4x)\] Like we have solved above \(\frac{d}{dx}(x^2+4x)=2x+4\) S our final answer would be \[\frac{1}{2}(x^2+4x)^{\large -\frac{1}{2}}*(2x+4)\]

OpenStudy (anonymous):

We still have to simplify our answer but read what I did and ask any question if you have any.